OpenStudy (anonymous):

Find the vertex, the equation of the axis of symmetry, and the y- intercept. y= x^2 - 9 Please show

pooja195 (pooja195):

The radius is 4.5/2 = 2.25 inches the area of the CD is A=πr2 A= 3.14*2.25*2.25= 15.896 sq in the inner hole has area 1.8 sq inches, which we subtract away: A= 15.896-1.8= 14.096 round to the nearest hundred A= 14.10